# LEGNet: Lightweight Edge-Gaussian Driven Network for Low-Quality Remote Sensing Image Object Detection

### v2

@@ 2.2 Backbone Networks for RSOD
To address this challenge, we propose LEGNet—a lightweight edge-Gaussian network tailored for low-quality RSOD. To our knowledge, LEGNet is the first RSOD backbone to integrate traditional image processing techniques (edge detection and Gaussian filtering) into a modern deep learning framework, offering a novel approach to robust feature extraction in degraded scenarios.
@@ 3.2 LoG-Stem Layer
The output of the LoG filter is activated $+$ normalized (AN), and added to the input via a residual connection: $\mathbf{F}_{LoG}=\texttt{Norm}\left(\mathbf{I}+\texttt{AN}\left(\texttt{LoG}^{7\times 7}_{1.0}\left(\texttt{Conv}^{7\times 7}_{2D}(\mathbf{I})\right)\right)\right).$ This residual-like design preserves image details and promotes stable gradient flow. $\mathbf{F}_{LoG}$ is then passed through two $3\times 3$ convolutional layers, with the second using stride 2 for downsampling: $\mathbf{F}_{1}=\texttt{ConvD}^{3\times 3}_{2D}\left(\texttt{Conv}^{3\times 3}_{2D}(\mathbf{F}_{LoG})\right),$ reducing spatial size to $\frac{H}{2}\times\frac{W}{2}$.
@@ 3.3.2 Gaussian Modeling
In Stages 2-4, edges are blurred with decreasing resolution, but object features tend to exhibit a Gaussian distribution centered around a specific point. To better capture these object features, Gaussian convolution kernels are employed, which help to smooth feature variations and suppress noise, thereby preserving critical object structures and preventing their degradation during the forward propagation of the network. As described in Eq. 2, the Gaussian filter assigns higher weights to pixels closer to the center of the kernel, emphasizing prominent features in the image while suppressing background noise. This is effective for stabilizing feature representations, enabling the network to focus on the most informative regions, and minimizing artifacts from high-frequency noise. We apply a fixed $5\times 5$ Gaussian kernel via depthwise convolution: $\mathbf{A}_{gauss}=G^{5\times 5}_{1.0}(\mathbf{F}_{in}),$ which smooths each channel, emphasizing central pixels and suppressing noise.
@@ 3.4 Macro Designs
Table 1: Performance comparison on the DOTA-v1.0 test set [36] under single-scale training and testing protocols. Results for other methods are sourced from [1, 25]. Red and blue indicate the best and second-best results per column, respectively.
@@ 4.1 Implementation Details
For training and evaluation, we followed the configurations from previous works [28, 19, 1, 25]. Specifically, for DOTA-v1.0, DOTA-v1.5, and DIOR-R, we applied single-scale training/testing with a scale factor of 3×3$\times$3 ×. DOTA images were cropped into $024\times 1,$ 024 × 1 , 024 patches with 200-pixel overlap, while DIOR-R used $800\times 800$ inputs. For FAIR1M-v1.0, we followed [19] and adopted multi-scale training/testing (scales: 0.5, 1.0, 1.5) with $024\times 1,$ 024 × 1 , 024 crops and 500-pixel overlap. For VisDrone2019, we used single-scale training/testing with a 1×1$\times$1 × scale and $1,333\times 800$ input size. The AdamW optimizer [22] was used with a momentum of 0.9 and weight decay of 0.05, an initial learning rate of 0.0002. A cosine learning rate schedule [23] with warm-up was adopted. Random resizing and flipping were applied during training. The test stage used the same resolution as training for consistency.
@@ 4.2 Quantitative Results
Table 3: Comparison of mAP and FPS with various backbone networks using the O-RCNN [37] detector on the DOTA-v1.0 test set. Inference speed, with detector, is measured on a single RTX 3090 GPU. Params (#P) and FLOPs are calculated for backbones only.
Method PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC CC mAP $\uparrow$
@@
Tab. 3 presents a comparative evaluation of LEGNet integrated with O-RCNN [37]. LEGNet-S achieves state-of-the-art performance with 80.03% mAP, outperforming all baselines while maintaining moderate complexity (12.7M params, 65.4G FLOPs) and a competitive speed (20.9 FPS), which is 1.77×1.$7$ faster than ARC-R50 [28].
@@ 4.2.4 Performance on DIOR-R
Table 6: Experimental results on the DIOR-R test set [4]. #P and FLOPs were tested by $800\times 800$.
@@ 4.3 Qualitative Results
Figure 3: Visualization of detection results on DOTA-v1.0 test set [36]. Input images resolution were $024\times 1,$ 024 × 1 , 024.
@@ 4.4 Ablation Studies
[…]
@@ 5 Conclusion
Appendix A Appendix
A.1 Datasets and Experimental Setup
@@
FAIR1M-v1.0 95,396 48,701 $>$1,000,000 5 (37 sub) 682$\times$682, 1,024$\times$1,024, 2,048$\times$2,048 Large-scale, fine-grained categories, geographic metadata, orientations
VisDrone2019 6,471 548 $>$2,600,000 10 480$\times$360 to 2,000$\times$1,500 (Varies) Drone-based, dense targets, complex backgrounds, diverse scenarios
Table 9: Comparison of Remote Sensing Object Detection Datasets
@@ 4 $\frac{H}{32}\times\frac{W}{32}$ DRFD Module $128/256$ $256/512$
[LEG Block] ×$\times N_{4}$ $/256$ 256 $/512$ 512
Number of Block [$N_{1}$, $N_{2}$, $N_{3}$, $N_{4}$] = [1, 4, 4, 2]
@@ A.2 LEGNet Configuration Details
Core Blocks: This stage includes $N_{1}$ repetitions of the ‘[LEG Block]’. Both Tiny and Small versions maintain the channel dimensions within these blocks (Tiny: $/32$ 32; Small: $/64$ 64), indicating a focus on learning hierarchical features without further channel changes at this sub-stage.
@@
Core Blocks: This stage comprises $N_{2}$ ‘[LEG Block]’ repetitions. Channels are maintained within these blocks (Tiny: $/64$ 64; Small: $/128$ 128), suggesting the primary role of these blocks is to refine features at the current resolution.
@@
Core Blocks: $N_{3}$ ‘[LEG Block]’ repetitions are used, preserving channel dimensions (Tiny: $/128$ 128; Small: $/256$ 256).
@@
Core Blocks: $N_{4}$ ‘[LEG Block]’ repetitions are present, maintaining channel dimensions (Tiny: $/256$ 256; Small: $/512$ 512).
@@
Block Repetitions. The number of ‘LEG Block’ repetitions for each stage, denoted as $[N_{1},N_{2},N_{3},N_{4}]$, is consistently set to $[1,4,4,2]$ for both Tiny and Small configurations.
A.3 Visualization Results
@@
The visual results on the DOTA-v1.0 test set are presented in Figs. 4, 5, 6, 7 and 8. We conducted a visual analysis using SOTA backbones, including ARC-R50 [28] and PKINet [1], both of which were designed specifically for RSOD tasks. The visualization results of ARC-R50 were based on the MMDetection toolbox, while ResNet-50 [12], PKINet-S and LEGNet-S visualization results were based on the MMRotate toolbox. All models were developed based on the O-RCNN [37] detector.
@@
Figure 4: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1024$\times$1024.
@@
Figure 5: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1024$\times$1024.
@@
Figure 6: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1024$\times$1024.
@@
Figure 7: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1024$\times$1024.
@@
Figure 8: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1024$\times$1024.

### v3

# LEGNet: A Lightweight Edge-Gaussian Network for Low-Quality Remote Sensing Image Object Detection

@@ 2.2 Backbone Networks for RSOD
To address this challenge, we propose LEGNet—a lightweight edge-Gaussian network tailored for low-quality RSOD. To the best of our knowledge, LEGNet is the first RSOD backbone to integrate traditional image processing techniques (edge detection and Gaussian filtering) into a modern deep learning framework, offering a novel approach to robust feature extraction in degraded scenarios.
@@ 3.2 LoG-Stem Layer
The output of the LoG filter is activated ++ normalized (AN), and added to the input via a residual connection: $\mathbf{F}_{LoG}=\texttt{Norm}\left(\mathbf{I}+\texttt{AN}\left(\texttt{LoG}^{7\times 7}_{1.0}\left(\texttt{Conv}^{7\times 7}_{2D}(\mathbf{I})\right)\right)\right).$ This residual-like design preserves image details and promotes stable gradient flow. $\mathbf{F}_{LoG}$ is then passed through two $3\times 3$ convolutional layers, with the second using stride 2 for downsampling: $\mathbf{F}_{1}=\texttt{ConvD}^{3\times 3}_{2D}\left(\texttt{Conv}^{3\times 3}_{2D}(\mathbf{F}_{LoG})\right),$ reducing spatial size to $\frac{H}{2}\times\frac{W}{2}$.
@@ 3.3.2 Gaussian Modeling
In Stages 2-4, edges are blurred with decreasing resolution, but object features tend to exhibit a Gaussian distribution centered around a specific point. To better capture these object features, Gaussian convolution kernels based on a Gaussian spatial prior are employed, which help to smooth feature variations and suppress noise, thereby preserving critical object structures and preventing their degradation during the forward propagation of the network. As described in Eq. 2, the Gaussian filter assigns higher weights to pixels closer to the center of the kernel, emphasizing prominent features in the image while suppressing background noise. This is effective for stabilizing feature representations, enabling the network to focus on the most informative regions, and minimizing artifacts from high-frequency noise. We apply a fixed $5\times 5$ Gaussian kernel via depthwise convolution: $\mathbf{A}_{gauss}=G^{5\times 5}_{1.0}(\mathbf{F}_{in}),$ which smooths each channel, emphasizing central pixels and suppressing noise.
@@ 3.4 Macro Designs
Table 1: Performance comparison on DOTA-v1.0 test set [36] under single-scale training and testing. Results for other methods are sourced from [1, 25]. “#P” is calculated for backbone + detector. Red and blue indicate the best and second-best results per column, respectively.
@@ 4.1 Implementation Details
[…]
@@ 4.2 Quantitative Results
Table 3: Comparison of mAP and FPS with various backbone networks using the O-RCNN [37] detector on the DOTA-v1.0 test set. Inference speed, with detector, is measured on a single RTX 3090 GPU. “#P” and FLOPs are calculated for backbones only.
Method PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC $C$ mAP $\uparrow$
@@
[…]
@@ 4.2.4 Performance on DIOR-R
Table 6: Experimental results on the DIOR-R test set [4]. #P and FLOPs were tested by 800 $\times$ 800 with backbone + detector.
@@ 4.3 Qualitative Results
Figure 3: Visualization of detection results on DOTA-v1.0 test set [36]. Input images resolution were 1,024 $\times$ 1,024.
@@ 4.4 Ablation Studies
LoG- -Stem EGA #P $\downarrow$ FLOPs $\downarrow$ mAP $\uparrow$ AP for each category ($\%$) $\uparrow$
@@ 5 Conclusion
Acknowledgements: This work was supported in part by NSFC Key Project of International (Regional) Cooperation and Exchanges (No. 61860206004), NSFC Key Project of Joint Fund for Enterprise Innovation and Development (No. U20B2068, U24A20342) and National Natural Science Foundation of China (No. 61976004).
Supplementary Materials for “LEGNet: A Lightweight Edge-Gaussian Network for Low-Quality Remote Sensing Image Object Detection”
This supplementary document provides additional details to complement the main paper. It aims to facilitate reproducibility and deepen the understanding of our proposed LEGNet, including its architectural design, core components, and extensive experimental validation.
The appendix is organized as follows:
  • Appendix A provides detailed descriptions of the datasets used for evaluation and the experimental setup.
  • Appendix B presents the specific architectural configurations for the different variants of LEGNet (Tiny and Small).
  • Appendix C includes an in-depth discussion on our key design choices, particularly the rationale for employing explicit priors and the sensitivity of hyperparameters.
  • Appendix D offers a visual analysis of intermediate feature responses, providing insight into how LEGNet enhances feature representations compared to other methods.
  • Appendix E showcases additional qualitative results, visualizing detection performance on the DOTA-v1.0 benchmark.
Appendix A Datasets and Experimental Setup
@@
FAIR1M-v1.0 95,396 48,701 >>1,000,000 5 (37 sub) 682 $\times$ 682, 1,024 $\times$ 1,024, 2,048 $\times$ 2,048 Large-scale, fine-grained categories, geographic metadata, orientations
VisDrone2019 6,471 548 >>2,600,000 10 480 $\times$ 360 to 2,000 $\times$ 1,500 (Varies) Drone-based, dense targets, complex backgrounds, diverse scenarios
Table 9: Comparison of Remote Sensing Object Detection Datasets
@@
Appendix B LEGNet Configuration Details
@@ 4 $\frac{H}{32}\times\frac{W}{32}$ DRFD Module $128/256$ $256/512$
[LEG Block] $\times N_{4}$ $256/256$ $512/512$
Number of Block $N_{1}$, $N_{2}$, $N_{3}$, $N_{4}$] = [1, 4, 4, 2]
@@
Core Blocks: This stage includes $N_{1}$ repetitions of the ‘[LEG Block]’. Both Tiny and Small versions maintain the channel dimensions within these blocks (Tiny: $32/32$; Small: $64/64$), indicating a focus on learning hierarchical features without further channel changes at this sub-stage.
@@
Core Blocks: This stage comprises $N_{2}$ ‘[LEG Block]’ repetitions. Channels are maintained within these blocks (Tiny: $64/64$; Small: $128/128$), suggesting the primary role of these blocks is to refine features at the current resolution.
@@
Core Blocks: $N_{3}$ ‘[LEG Block]’ repetitions are used, preserving channel dimensions (Tiny: $128/128$; Small: $256/256$).
@@
Core Blocks: $N_{4}$ ‘[LEG Block]’ repetitions are present, maintaining channel dimensions (Tiny: $256/256$; Small: $512/512$).
@@
Block Repetitions. The number of ‘LEG Block’ repetitions for each stage, denoted as $N_{1},N_{2},N_{3},N_{4}]$, is consistently set to $[1,4,4,2]$ for both Tiny and Small configurations.
Appendix C Discussion on Design Choices
This section elaborates on key design choices within LEGNet, providing the rationale for embedding explicit priors over purely end-to-end learning and discussing the sensitivity of crucial hyperparameters.
C.1 Rationale for Employing Explicit Priors
A natural question regarding our approach is why we chose to explicitly encode priors, such as edge and Gaussian features, rather than relying on a sufficiently deep or complex network (e.g., a Transformer-based model) to learn them automatically. While modern deep networks possess immense learning capabilities, our approach of embedding explicit priors offers several distinct advantages, particularly in the context of our goal to build a lightweight and robust backbone for RSOD.
  • Data and Parameter Efficiency: Learning fundamental concepts like edges or Gaussian-like attention from scratch is a parameter-intensive task. By embedding these priors through parameter-free operators (our LoG-Stem and the fixed Scharr/Gaussian filters in the EGA module), we provide the network with a strong, built-in ”head start.” This makes the model significantly more data-efficient and allows it to achieve high performance with a much lower parameter count—a core objective of our lightweight LEGNet design.
  • Robustness to Degradation: This is a cornerstone of our motivation. End-to-end models learn features based on statistical patterns in the training data. When input images are degraded (e.g., due to blur, low contrast, or noise), these patterns can become weak or distorted, causing learned filters to fail. In contrast, our explicit edge detectors are deterministic operators that can reliably extract structural information even from degraded signals. By providing the network with this robust edge map, we ensure that subsequent layers receive meaningful structural cues, enhancing the model’s resilience to poor imaging conditions.
  • Guided Learning and Regularization: Explicit priors act as a strong inductive bias, guiding the network to focus on structurally relevant information from the earliest stages. This serves as a form of regularization, discouraging the model from overfitting to spurious textures or background noise. As discussed in our macro design (Sec. 4.5), this biases the model towards learning ”the right features for the right reasons,” leading to better generalization.
  • Interpretability: The use of well-defined operators like Laplacian of Gaussian provides a degree of interpretability to the early-stage feature extraction process. We know precisely that the initial layers are enhancing edges, which aligns with human visual processing and provides clearer insight into the model’s behavior, as supported by our feature visualizations in Fig. 4 of the Appendix.
In summary, our choice is not based on the premise that deep networks *cannot* learn these features, but rather that explicitly encoding them provides a more efficient, robust, and targeted pathway to building a high-performing lightweight model for the specific challenges of RSOD.
C.2 Discussion on Hyperparameter Sensitivity
The hyperparameters of the Gaussian kernel, namely its size and bandwidth (standard deviation, $\sigma$), are influential components of the EGA module. While a full ablation study was beyond the scope of our primary investigation, we discuss their roles and the reasoning behind our choices here.
  • Gaussian Kernel Size: The kernel size determines the spatial extent of the feature aggregation. A larger kernel incorporates context from a wider neighborhood, beneficial for larger objects, but risks over-smoothing details and blurring together small, dense objects. Conversely, a smaller kernel preserves fine details but may fail to capture sufficient context. Our selected kernel size was determined during preliminary experiments to strike a balance suitable for the multi-scale nature of objects in datasets like DOTA and FAIR1M. The chosen value proved effective at capturing salient features without significant information loss.
  • Gaussian Bandwidth ($\sigma$): The bandwidth controls the decay rate of the Gaussian function. A small $\sigma$ creates a sharp filter that heavily prioritizes the central feature, while a large $\sigma$ creates a smoother filter that gives more uniform weight to the neighborhood. The value of $\sigma$ is often set proportionally to the kernel size to ensure the distribution fits naturally within the kernel’s window. Our implementation follows this standard practice, a well-established heuristic in computer vision for ensuring stability.
Crucially, while these parameters are important, we found the model’s performance to be robust to minor variations around our chosen values. This robustness is largely because the subsequent learnable layers (e.g., the 1x1 convolutions within our LEG Blocks) grant the network the flexibility to adapt and recalibrate the features generated by the fixed-parameter Gaussian module. Therefore, the overall architecture is not hyper-sensitive to these specific values, and the selected parameters represent a stable and effective configuration for the task.
Appendix D Analysis of Intermediate Feature Responses
To provide deeper insight into LEGNet’s mechanism, we visualize the intermediate feature maps. As shown in Fig. 4, we compare the feature responses of our LEGNet-S with a strong competitor, PKINet-S, at different stages of the backbone: the initial stem layer, the output of Stage 1, and the output of Stage 2.
The visualizations reveal distinct differences in feature representation. In the stem layer and Stage 1, LEGNet (bottom row) demonstrates a superior ability to extract comprehensive and complete edge information across the entire image compared to PKINet (top row). This is attributed to our LoG-Stem layer and the EGA module, which explicitly enhance structural details from the outset. As the network deepens into Stage 2, LEGNet exhibits a more focused attention on salient object features, effectively suppressing background noise and highlighting regions of interest. In contrast, the feature responses from PKINet appear more diffuse. This comparison visually substantiates our claim that by integrating explicit edge and Gaussian priors, LEGNet learns more robust and meaningful feature representations, particularly in the crucial early stages of processing. Due to the highly abstract nature of features in deeper stages (3 and 4), we focus on these initial stages where the impact of our design is most visually interpretable.
Figure 4: Visualization of intermediate feature maps from PKINet-S (top row) and our LEGNet-S (bottom row). LEGNet’s early stages capture more complete edge details, while Stage 2 demonstrates enhanced focus on salient object regions. This supports the effectiveness of our proposed design in refining feature representations for robust detection.
Appendix E Visualization Results
@@
The visual results on the DOTA-v1.0 test set are presented in Figs. 5, 6, 7, 8 and 9. We conducted a visual analysis using SOTA backbones, including ARC-R50 [28] and PKINet [1], both of which were designed specifically for RSOD tasks. The visualization results of ARC-R50 were based on the MMDetection toolbox, while ResNet-50 [12], PKINet-S and LEGNet-S visualization results were based on the MMRotate toolbox. All models were developed based on the O-RCNN [37] detector.
@@
Figure 5: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1,024 $\times$ 1,024.
@@
Figure 6: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1,024 $\times$ 1,024.
@@
Figure 7: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1,024 $\times$ 1,024.
@@
Figure 8: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1,024 $\times$ 1,024.
@@
Figure 9: Visualization of detection results on DOTA-v1.0 test set. Input images resolution were 1,024 $\times$ 1,024.